OpenStudy (anonymous):

A company that manufactures storage bins for grains made a drawing of a silo. The silo has a conical base, as shown below. What is the total volume of grains that can be stored in the silo? Answer 14π ft3 18π ft3 24π ft3 30π ft3

OpenStudy (oaktree):

As shown below, please!

OpenStudy (oaktree):

Split it into a cylinder and a cone.

OpenStudy (oaktree):

You should be able to take it from there.

OpenStudy (anonymous):

you would think... but im soo tired can you give me the equation i can def. do it from there

OpenStudy (anonymous):

For the volume of a cylinder multiply the area of the circle \(\pi r^2\) times the height.

OpenStudy (anonymous):

For the cone \(\dfrac{1}{3} \pi r^2 h\)

OpenStudy (anonymous):

ok so 12.56 is the volume of the cylinder

OpenStudy (anonymous):

But you have to wait until calculus to find out where the \(\dfrac{1}{3}\) comes from.

OpenStudy (anonymous):

mmk and the volume of the cone would be 39.56 or 40 right and then i add the two?

OpenStudy (anonymous):

Your choices all have\(\pi\) in them so you may want to not simplify your answer.

OpenStudy (anonymous):

so i would add them up and then like devide by pi to simplify?

OpenStudy (anonymous):

Not exactly, the area of the cylinder is \(\pi(2)^2(6)=24\pi\). You want to leave it as \(24\pi\)

OpenStudy (anonymous):

ok got it so i just leave pi out of the equation since its in my answer

OpenStudy (anonymous):

Yes you got it.

OpenStudy (anonymous):

thank-you so much i know im tough to help but im not good at math and have been working since 6 this morning your help was greatly appreciated!!!!